OpenStudy (anonymous):

a circle has a circumference of 10 cm The radius of this circle is closest to A. 1.3cm b.1.6cm c.1.8cm d.3.2cm e.5.0cm a circle has a circumference of 10 cm The radius of this circle is closest to A. 1.3cm b.1.6cm c.1.8cm d.3.2cm e.5.0cm @Mathematics

OpenStudy (asnaseer):

use the formulae that relates the circumference C to its radius r:\[C=2\pi r\]to get:\[r=C/2\pi\]

OpenStudy (asnaseer):

you can approximate pi as:\[\pi\approx \frac{22}{7}\]

OpenStudy (anonymous):

how, what answer did u get

OpenStudy (asnaseer):

so then you get:\[r\approx \frac{C}{2*\frac{22}{7}}\]\[r\approx \frac{C}{\frac{44}{7}}\]\[r\approx C*\frac{7}{44}\]